A heat exchanger is being installed as part of a plant modernization program. It costs $80,000, including




installation, and is expected to reduce the overall plant fuel cost by $20,000 per year. Estimates of the




useful life of the heat exchanger range from an optimistic 12 years to a pessimistic 4 years. The most




likely value is 5 years. Assume the heat exchanger has no salvage value at the end of its useful life.




Determine the pessimistic, most likely, and optimistic rates of return.

1
Expert's answer
2021-12-02T10:38:34-0500

The pessimistic rate of return is:

"R = (20,000\u00d7 - 80,000)\/80,000\u00d7100" % = 0%.

The most likely rate of return is:

"R = (20,000\u00d75 - 80,000)\/80,000\u00d7100" % = 25%.

The optimistic rate of return is:

"R = (20,000\u00d712 - 80,000)\/80,000\u00d7100" % = 200%.